# Otter-Knowledge: benchmarks of multimodal knowledge graph representation learning from different sources for drug discovery

### v1

@@ Abstract
Recent research in representation learning utilizes large databases of proteins or molecules to acquire knowledge of drug and protein structures through unsupervised learning techniques. These pre-trained representations have proven to significantly enhance the accuracy of subsequent tasks, such as predicting the affinity between drugs and target proteins. In this study, we demonstrate that by incorporating knowledge graphs from diverse sources and modalities into the sequences or SMILES representation, we can further enrich the representation and achieve state-of-the-art results on established benchmark datasets. We provide preprocessed and integrated data obtained from 7 public sources, which encompass over 30M triples. Additionally, we make available the pre-trained models based on this data, along with the reported outcomes of their performance on three widely-used benchmark datasets for drug-target binding affinity prediction found in the Therapeutic Data Commons (TDC) benchmarks. Additionally, we make the source code for training models on benchmark datasets publicly available. Our objective in releasing these pre-trained models, accompanied by clean data for model pretraining and benchmark results, is to encourage research in knowledge-enhanced representation learning.
@@ 1 Introduction
Developing a concise representation of proteins and small molecules is a crucial task in AI-based drug discovery. Recent studies [22, 26] have focused on utilizing large databases of protein sequences or molecules for self-supervised representation learning. These representations are then fine-tuned using limited labeled data for tasks like predicting the binding affinity between drugs and targets. In the field of protein representation learning, [38] and [39] have demonstrated that enhancing protein representations with additional information from knowledge graphs, such as comprehensive textual data from the gene ontology [1], can enhance the performance of pre-trained representations on various protein properties and protein-protein interaction tasks.
While early results concerning knowledge-enriched representations for proteins show promise, there is a notable absence of openly available, carefully curated, integrated, and readily usable datasets in the research community for studying pre-training methods. This scarcity motivates us to preprocess datasets from diverse knowledge sources, incorporating abundant factual information and modalities, and make the data available alongside the pre-trained models, along with the prediction accuracy of downstream models. Our primary objective is to establish foundational datasets for research on multimodal knowledge-enhanced representation learning and evaluate the outcomes against established benchmarks for downstream tasks. In addition to presenting new state-of-the-art findings on standard benchmark datasets, we offer a comprehensive discussion that shares the insights gained during the creation of pretrained models, as well as the research challenges that must be overcome to develop knowledge-enhanced foundation models for therapeutic sciences. The preprocessed datasets, pretrained models and the code for benchmark models training are open-sourced and available in our project github repository [^1] https://github.com/IBM/otter-knowledge.
@@ 2 Multimodal knowledge representation learning
The diagram in Figure 1 illustrates the overall process of our system named Otter-Knowledge. This process involves constructing multimodal knowledge graphs from diverse sources, generating initial embeddings for each modality using pretrained models available in the model zoo, and subsequently improving the embeddings by incorporating information from the knowledge graphs through the utilization of graph neural networks (GNN).
@@ 2.1 Multimodal Knowledge graph construction
A Multimodal Knowledge Graph (MKG) is a directed labeled graph where labels have well-defined meanings, and each node has a modality, a particular mode that qualifies its type (text, image, protein, molecule, etc.). We consider two node subsets: entity nodes (or entities), which corresponds to concepts in the knowledge graph (for example protein, or molecule), and attribute nodes (or attributes), which represent qualifying attributes of an entity (for example the mass of a molecule, or the description of a protein). We refer to an edge that connect an entity to an attribute as data property, and an edge that connect two entities as object property. Each node in the graph has a unique identifier, and a unique modality (specified as a string).
We developed a framework for automating the construction of a multimodal knowledge graph by extracting and fusing data from a variety of sources, including text delimited files, JSON, and proprietary data sources [32]. The framework takes as input a schema file (specified in JSON), which declaratively describe how to build the desired graph from a set of data sources.
The framework that builds the MKG ensures that each triple is unique, and it automatically merges entities having the same unique identifier, but whose data is extracted from different data sources. It is also possible to use the special relation sameAs  [^2] We borrow the semantic of owl:sameAs - see https://www.w3.org/TR/owl-ref/#sameAs-def to indicate that two entities having different unique identifiers are to be considered as the same entity. The sameAs relation is useful when creating a MKG from multiple partially overlapping data sources; when the graph is built. Additionally, it is possible to build an MKG incrementally, by merging two or more graphs built using different schemas; the merge operation automatically fuses entities having the same unique identifier or the same value of a distinctive attribute (for example the same sequence for two proteins).
The framework builds the graph in memory, but can also provide support for building the graph using a database on disk; the graph triples can also be serialised using GML  [^3] https://en.wikipedia.org/wiki/Graph_Modelling_Language or any RDF  [^4] https://www.w3.org/RDF/ serialization formats. Finally, the framework provides scalable functionalities for parallel and GPU-based computation of the initial embedding vectors for the nodes; such embeddings are computed based on the modality of each node.
@@ 2.2 Computing initial embeddings
As explained before, the MKG contains nodes representing entities and nodes representing attributes of those entities. In the MKG, each node has a modality assigned, e.g: entity nodes can have a modality Protein or Drug, nodes containing text could have a modality text. We assign a model for each one of the modalities in our graph, as specified by the user in the schema. The models, which we refer to as handlers, are capable of preprocessing the values in the nodes and computing their initial embeddings. Our framework allows to easily retrieve all the nodes in the graph with the same modality to efficiently compute the initial embeddings with the assigned handler, facilitating parallelization, GPU utilization, batching, and avoiding the need to load different models in memory simultaneously.
Some of these modalities do not have a handler assigned, like Protein, and therefore, no initial embedding is computed for them. For each modality, only one handler is being used, although it is possible to change the handler assigned. For instance, for SMILES it is possible to use morgan-fingerprint or MolFormer. These are the handlers that we have used for computing the initial embeddings of the graph:
  • morgan-fingerprint We use the Morgan fingerprint in RDKit [^5] https://www.rdkit.org/docs/index.html for processing the SMILES. Specifically, we use GetMorganFingerprintAsBitVect. We use by default a shape (also known as nBits or size) of 2048, and a radius of 2. If RDKit is not able to get the Morgan fingerprint of the SMILES we return an embedding of the same shape full of zeros.
@@
  • protein-sequence-mean For the protein sequences, we use esm1b_t33_650M_UR50S [26]. We use the 33rd layer as representation layer, and compute the mean for the contacts. For the batch_converter we use $truncation\_seq\_length=1022$.
  • text For the textual values, we use the ‘sentence-transformers/paraphrase-albert-small-v2‘ [24] model available in the Huggingface Hub [^6] https://huggingface.co/sentence-transformers/paraphrase-albert-small-v2.
@@ 2.3 Pretraining with inductive R-GNN
To improve representations of MKG entities we assimilate initial embeddings with the information about the MKG structure, i.e., connection patterns between graph nodes. To this end we train a Graph Neural Network (GNN) [40] that propagates initial embeddings through a set of layers that upgrade input embedding according to the node neighbours. The architecture of GNN consists of two main blocks: encoder and decoder. For encoder we first define a projection layer which consists of a set of linear transformations for each node modality and projects nodes into common dimensionality, then we apply several multi-relational graph convolutional layers (R-GCN) [29] which distinguish between different types of edges between source and target nodes by having a set of trainable parameters for each edge type. For decoder we consider link prediction task, which consists of a scoring function that maps each triple of source and target nodes and the corresponding edge and maps that to a scalar number defined over interval $[0;1]$. Decoder and scoring function, and details about the hyperparameters are reported in the Appendix.
@@
For link prediction, we consider three choices of scoring functions: DistMult, TransE and a Binary Classifier that are commonly used in the literature. The outcomes of scoring of each triple are then compared against actual labels using negative log likelihood loss function. Certain pretraining datasets, like Uniprot, contain numerical data properties. Hence, we incorporate an extra regression objective aimed at minimizing the root mean square error (MSE) of the predicted numerical data properties. In the learning process, we combine the regression objective and the link prediction objective to create a single objective function.
@@
We note that to train GNN with link prediction we need to provide network with both positive and negative examples. While positive examples come from the MKG itself, negative triples are generated synthetically. For this for each relation type we extract a set of admissible source and target nodes and then we randomly sample source and target from the corresponding admissible sets. We use an equal ratio between positive and negative links in the pretraining.
@@
Due to the integration of data from various sources, the size of the integrated data can become quite large. For instance, the combination of Uniprot, ChemBL, and BindingDB necessitates over 200GB of CPU memory for training. However, training a graph neural network (GNN) on such a massive graph using limited GPU memory is highly inefficient. To address this, we employ a graph auto-scaling approach (GAS) described in the reference [8]. This method divides the graph into smaller partitions using Metis [^7] https://github.com/KarypisLab/METIS. GAS performs training on each partition separately, so it is able scale to arbitrarily large graphs. To avoid information loss due to the connection between partitions, it keeps node embeddings of previous training step in CPU memory thus saving GPU memory during training and uses this historical embeddings to update the nodes inside a partition. Please refer to Appendix for details about the hyper-parameter settings of GAS.
Information flow control and noisy link prediction
One crucial aspect of pretraining the GNN involves addressing the disparity between the data accessible during pretraining and the data accessible during subsequent tasks. Specifically, during pretraining, there are numerous attributes associated with proteins or drugs, whereas during downstream fine-tuning, only amino acid sequences and SMILES are available. Consequently, during pretraining, we explore two scenarios: one which controls the information propagated to the Drug/Protein entities and one without such control. In our experiments, we present results for both cases to provide an insight on the impact of restricting information flow during pretraining on the subsequent tasks. An additional significant consideration is the presence of noisy links within the up-stream data and how they affect the downstream tasks. To investigate the potential impact on these tasks, we manually handpick a subset of links from each database that are relevant to drug discovery (see details in the Appendix). We then compare the outcomes when training the GNN using only these restricted links versus using all possible links present in the graphs.
@@ 3 Pretraining datasets collections
Table 1 summarises the pretrained datasets. In this subsection, we discuss data preprocessing methods. For details about datasets and schema, readers can refer to the appendices and the github repository.
@@
UBC 6,207,654 Pro teins/Drugs sequences, SMILES text, number, category Open Yes
PrimeKG 12,757,257 Proteins/Drugs/ Diseases sequences, SMILES text Open Yes
DUDe 40,216 Proteins/Drugs sequences, SMILES Open Yes
STITCH 12,621,873 Proteins/Chemicals sequences, SMILES Open Yes
@@
[5] comprises of 573,227 proteins from SwissProt, which is the subset of manually curated entries within UniProt, including attributes with different modalities like the sequence (567,483 of them), full name, organism, protein family, description of its function, catalytics activity, pathways and its length. The number of edges are 38,665 of type target_of from Uniprot ids to both ChEMBL and Drugbank ids, and 196,133 interactants between Uniprot protein ids.
@@
[16] consists of 2,656,221 data points, involving 1.2 million compounds and 9,000 targets. Instead of utilizing the affinity score, we generate a triple for each combination of drugs and proteins. In order to prevent any data leakage, we eliminate overlapping triples with the TDC DTI dataset. As a result, the dataset concludes with a total of 2,232,392 triples.
@@
[9] comprises of drug-like bioactive molecules, 10,261 ChEMBL ids with their corresponding SMILES were downloaded from OpenTargets [18], from which 7,610 have a sameAs link to drugbank id molecules.
@@
[13] comprises of detailed chemical data on 9,749 drugs (such as SMILES, description, indication, mechanism of action, affected organism, average mass, toxicity, calculated and experimental properties, absorption), classification, drug pathways and 1,301,422 drug interactions
@@
[17] comprises a collection of 22,886 active compounds and their corresponding affinities towards 102 targets. For our study, we utilized a preprocessed version of the DUDe [31], which includes 1,452,568 instances of drug-target interactions. To prevent any data leakage, we eliminated the negative interactions and the overlapping triples with the TDC DTI dataset. As a result, we were left with a total of 40,216 drug-target interaction pairs.
@@
(the Precision Medicine Knowledge Graph) [4] integrates 20 biomedical resources, it describes 17,080 diseases with 4 million relationships. PrimeKG includes nodes describing Gene/Proteins (29,786) and Drugs (7,957 nodes). The MKG that we built from PrimeKG contains 13 modalities, 12,757,300 edges (154,130 data properties, and 12,603,170 object properties), including 642,150 edges describing interactions between proteins, 25,653 edges describing drug-protein interactions, and 2,672,628 describing interactions between drugs.
@@
(Search Tool for Interacting Chemicals) [34] is a database of known and predicted interactions between chemicals represented by SMILES strings and proteins whose sequences are taken from STRING database [33]. Those interactions are obtained from computational prediction, from knowledge transfer between organisms, and from interactions aggregated from other (primary) databases. For the MKG curation we filtered only the interaction with highest confidence, i.e., the one which is higher 0.9. This resulted into 10,717,791 triples for 17,572 different chemicals and 1,886,496 different proteins. Furthermore, the graph was split into 5 roughly same size subgraphs and GNN was trained sequentially on each of them by upgrading the model trained using the previous subgraph.

### v3

@@ Abstract
Recent research on predicting the binding affinity between drug molecules and proteins use representations learned, through unsupervised learning techniques, from large databases of molecule SMILES and protein sequences. While these representations have significantly enhanced the predictions, they are usually based on a limited set of modalities, and they do not exploit available knowledge about existing relations among molecules and proteins. In this study, we demonstrate that by incorporating knowledge graphs from diverse sources and modalities into the sequences or SMILES representation, we can further enrich the representation and achieve state-of-the-art results for drug-target binding affinity prediction in the established Therapeutic Data Commons (TDC) benchmarks. We release a set of multimodal knowledge graphs, integrating data from seven public data sources, and containing over 30 million triples. Our intention is to foster additional research to explore how multimodal knowledge enhanced protein/molecule embeddings can improve prediction tasks, including prediction of binding affinity. We also release some pretrained models learned from our multimodal knowledge graphs, along with source code for running standard benchmark tasks for prediction of biding affinity.
@@ 1 Introduction
Developing a concise representation of proteins and small molecules is a crucial task in AI-based drug discovery. Recent studies [22, 26] have focused on utilizing large databases of protein sequences or molecules for self-supervised representation learning. These representations are then fine-tuned using limited labeled data for tasks like predicting the binding affinity between drugs and targets. In the field of protein representation learning, [38] and [39] have demonstrated that enhancing protein representations with additional information from knowledge graphs (KGs), such as comprehensive textual data from the gene ontology (GO) [1], can enhance the performance of pre-trained representations on various protein properties and protein-protein interaction tasks. Despite promising initial findings, important research directions remain unexplored.
Firstly, it is important to investigate if joint robust representations of proteins, molecules and other associated entities (such as diseases) may improve downstream prediction tasks. A joint representation of each of these entities should take into account not only the intrinsic properties of the entity itself, but also its relations with other entities. In contrast to previous studies [38, 39], concentrating on protein sequences, GO terms, and text descriptions, our work encompasses a wider array of modalities—text, numbers, protein sequences, SMILES, categories, and entities like drugs, diseases, and pathways.
Secondly, previous studies [38, 39] assumed that the graphs or datasets used for training were carefully combined into a single source. However, in our research, we deal with different KGs built using several datasets from various sources. It’s worth highlighting that merging these datasets into a single source is complex due to the difficulty of aligning their structures automatically. For example, in the STITCH knowledge graph [34], the connection labeled “interaction with", denoting the relationship between chemicals and proteins, is established by analyzing their co-occurrence within a Pubmed abstract. This form of association might bear resemblance to the “target of" relation found in the Uniprot knowledge graph, however confirming their equivalence presents a challenging task because co-orcurrence in a Pubmed abstract does not mean a protein is a target of a drug. Moreover, creating a comprehensive graph by combining multiple ones demands significant computational resources. To address these issues, we propose an approach based on ensemble methods, and we show that it can effectively capture information from a subset of our multimodal knowledge graphs without merging them.
Finally, while there are datasets for studying the problem of drugs-proteins interactions, there are no publicly available multimodal knowledge graphs for this purpose. We aim to provide the research community with such graphs, by selectively integrating knowledge from existing datasets.
The contributions of this work are summarised as follows:
  • We release to the research community a set of multimodal knowledge graphs, and a set of models pretrained on those graphs using graph neural networks to derive protein and drug representations (https://github.com/IBM/otter-knowledge). We hope that this will foster further research in the area.
  • We provide experimental results showing that a multimodal knowledge enhanced representation surpasses the state of the art for predicting drugs-proteins interactions, even for challenging discovery scenario, where majority of entities in the test data (e.g., molecules) are unseen in the training data and might only have one available modality (e.g, its SMILES). Previous work focused on enhancing Protein Language Models (PLMs) with textual descriptions, and did not exploit KGs that capture heterogenous relations between entities and multimodal entity attributes (including molecules SMILES, proteins sequences, textual, numerical and categorical attributes).
  • We tackle the problem of learning from partially connected multimodal knowledge graphs via ensemble methods, yielding promising outcomes. We also study how our approach remains robust across diverse pretraining objectives, such as regression and link prediction.
@@ 2 Multimodal knowledge representation learning
The diagram in Figure 1 illustrates the overall process of our system named Otter-Knowledge. This process involves constructing multimodal knowledge graphs from diverse sources, generating initial embeddings for each modality using pretrained models available in the model zoo, and subsequently improving the embeddings by incorporating information from the knowledge graphs through the utilization of graph neural networks (GNN). We discuss each component of the given system in the following subsections.
@@ 2.1 Multimodal Knowledge graph construction
A Multimodal Knowledge Graph (MKG) is a directed labeled graph where labels for nodes and edges have well-defined meanings, and each node has a modality, a particular mode that qualifies its type (text, image, protein, molecule, etc.). We consider two node subsets: entity nodes (or entities), which corresponds to concepts in the knowledge graph (for example protein, or molecule), and attribute nodes (or attributes), which represent qualifying attributes of an entity (for example the mass of a molecule, or the description of a protein). We refer to an edge that connects an entity to an attribute as data property, and an edge that connects two entities as object property. Each node in the graph has a unique identifier, and a unique modality (specified as a string).
We developed a framework for automating the construction of a multimodal knowledge graph by extracting and fusing data from a variety of sources, including text delimited files, JSON, and proprietary data sources [32]. The framework takes as input a schema file (specified in JSON), which declaratively describes how to build the desired graph from a set of data sources; a user may extend predefined schema files used to create our multimodal KGs (see our github repository), or define custom ones.
The framework that builds the MKG ensures that each triple is unique, and it automatically merges entities having the same unique identifier, but whose data is extracted from different data sources. It is also possible to use the special relation sameAs  [^1] We borrow the semantic of owl:sameAs - see https://www.w3.org/TR/owl-ref/#sameAs-def to indicate that two entities having different unique identifiers are to be considered as the same entity. The sameAs relation is useful when creating a MKG from multiple partially overlapping data sources; when the graph is built. Additionally, it is possible to build an MKG incrementally, by merging two or more graphs built using different schemas; The merge operation automatically combines entities with matching unique identifiers or distinct attributes (e.g., proteins with the same sequence).
The framework builds the graph in memory, but can also provide support for building the graph using a database on disk; the graph triples can also be serialised using GML  [^2] https://en.wikipedia.org/wiki/Graph_Modelling_Language or any RDF  [^3] https://www.w3.org/RDF/ serialization formats. Finally, the framework offers scalable parallel and GPU-based computation of initial node embeddings, tailored to each node’s modality.
@@ 2.2 Computing initial embeddings
As explained before, the MKG contains nodes representing entities and nodes representing attributes of those entities. In the MKG, each node has a modality assigned, e.g: entity nodes can have a modality Protein or Drug, nodes containing text could have a modality text or protein sequence. We assign a model for each one of the modalities in our graphs, as specified by the user in the schema. The models, referred to as handlers, are capable of preprocessing the values in the nodes and computing their initial embeddings. Our framework allows to easily retrieve all the nodes in the graph with the same modality to efficiently compute the initial embeddings with the assigned handler, facilitating parallelization, GPU utilization, batching, and avoiding the need to load different models in memory simultaneously.
Handlers are exclusively assigned to nodes representing attributes. For each modality, only one handler is being used, although it is possible to change the handler assigned. For instance, for SMILES it is possible to use morgan-fingerprint or MolFormer. The comparison between different models is not within the scope of this work. These are the handlers that we have used for computing the initial embeddings of the graph:
  • morgan-fingerprint We use the Morgan fingerprint in RDKit [^4] https://www.rdkit.org/docs/index.html for processing the SMILES. We employ a shape of 2048 (nBits) and a radius of 2 for the Morgan fingerprint. If RDKit fails to generate the fingerprint, we supply an embedding of the same shape filled with zeros.
@@
  • protein-sequence-mean For the protein sequences, we use the mean token embeddings extracted from the 33rd layer of the esm1b_t33_650M_UR50S model [26]. All sequences are truncted to 1022 amino acids.
  • text For the textual values, we use the ‘sentence-transformers/paraphrase-albert-small-v2‘ [24]  [^5] https://huggingface.co/sentence-transformers/paraphrase-albert-small-v2.
@@
  • categorical, protein and molecule entities modalities do not possess handlers, resulting in the absence of an initial embedding. Initial embeddings start as zeros and improve through inductive GNN training, based on the neighborhood connections to their data properties and related entities. To illustrate this, an effective representation for an entity Protein is learned through iterative aggregation of information from adjacent nodes representing multimodal attributes (such as its sequence, other textual attributes like label, family, function) initialised with their respective handlers and interactions with other entity nodes (proteins, molecules or categories), provided such modalities and relations exist.
Using distinct models lets us leverage existing competitive pretrained language models for meaningful modality representations (e.g., gauging embedding similarity). This aids enhancing generalization, the GNN can exploit cross-modal links through KG geometry and it reduces the need for a larger training dataset to generalise to unseen entities, which would have been necessary if we had opted for a single unified model setup for all modalities.
@@ 2.3 Pretraining with inductive R-GNN
To improve representations of MKG entities we assimilate initial embeddings with the information about the MKG structure, i.e., connection patterns between graph nodes. To this end we train GNN [40] that propagates initial embeddings of attribute nodes through a set of layers that transform input embedding according to the entity data and object properties. Our GNN is inductive since it may compute entities embeddings only from their available neighbours and the corresponding initial embeddings. This allows embeddings computation of unseen nodes during the training.
The architecture of GNN adheres to a standard design and consists of two main blocks: encoder and decoder. For the encoder, we first define a projection layer, which consists of a set of linear transformations for each attribute node modality and projects those into a common dimensionality. Then, we apply several multi-relational graph convolutional layers (R-GCN) [29], which distinguish between different types of edges connecting source and target nodes by having a set of trainable parameters for each edge type. For the decoder, we consider a link prediction task, which consists of a scoring function that maps each triple of source and target nodes, along with their corresponding edge, to a scalar number defined over interval [0;1]01[0;1][ 0 ; 1 ].
@@
For link prediction, we evaluate three common scoring functions: DistMult, TransE, and a Binary Classifier. We compare their scores against actual labels using negative log likelihood loss. Additionally, for numerical attributes, we include a regression objective to minimize root mean square error (MSE) of predicted numerical properties. The learning process combines the link prediction and regression objectives into a single function. GNN hyperparameters values are detailed in the Appendix.
@@
To train the GNN for link prediction we need to provide both positive and negative examples. While positive examples come from the MKG itself, negative triples are generated synthetically. To achieve this, for each relation type we extract a set of admissible source and target nodes, subsequently, we randomly sample sources and targets from the corresponding admissible sets. We use an equal ratio between positive and negative links.
@@
Due to the integration of data from various sources, the size of the integrated data can become significantly large. For instance, the combination of Uniprot, ChemBL, and BindingDB requires over 200GB of CPU memory for training, which renders GNN training on such a large graph using limited GPU memory highly inefficient. To address this, we employ a graph auto-scaling approach (GAS) described in [8]. This method divides the graph into smaller partitions using Metis [^6] https://github.com/KarypisLab/METIS. GAS performs training on each partition separately, so it is able scale to arbitrarily large graphs. To avoid information loss due to the connection between partitions, it keeps node embeddings of previous training step in CPU memory thus saving GPU memory and uses this historical embeddings to update the nodes inside a partition. Please refer to Appendix for details about the hyper-parameter settings of Metis.
Information flow control
One crucial aspect of pretraining the GNN involves addressing the disparity between the data accessible during pretraining and the data accessible during subsequent tasks. Specifically, during pretraining, there are numerous attributes associated with proteins or drugs, whereas during downstream fine-tuning, only protein sequences and SMILES are available. Consequently, during pretraining, we explore two scenarios: one which controls the information propagated to the Drug/Protein entities and one without such control. Our experiments cover both cases to shed light on how constraining information flow during pretraining affects subsequent tasks.
Noisy link prediction
An additional significant consideration is the presence of noisy links within the up-stream data and how they affect the downstream tasks. To investigate the potential impact on these tasks, we manually handpick a subset of links from each database that are relevant to drug discovery (see details in the Appendix). We then compare the outcomes when training the GNN using only these restricted links versus using all possible links that are potentially noisy with respect to the downstream tasks.
@@ 3 Pretraining datasets collections
Table 1 summarizes the multimodal KGs used for pretraining our models; they are based on existing KGs and datasets, and in this section we discuss the data preprocessing steps that we applied to the original KGs and datasets. We open-sourced the preprocessed KGs: we refer to the appendices and our github repository for further details.
@@
Otter-UBC 6,207,654 Proteins/Drugs sequences, SMILES, text, number Open Yes
Otter-PrimeKG 12,757,257 Proteins/Drugs/Diseases sequences, SMILES, text Open Yes
Otter-DUDe 40,216 Proteins/Drugs sequences, SMILES Open Yes
Otter-STITCH 12,621,873 Proteins/Chemicals sequences, SMILES Open Yes
@@
[5] comprises 573,227 SwissProt proteins (from curated UniProt subset). The Otter-UBC KG combines all UniProt proteins with diverse attributes, including sequence (567,483 entries), full name, organism, protein family, function, catalytic activity, pathways and length. The KG also features 38,665 target_of edges linking UniProt IDs to ChEMBL and Drugbank IDs, along with 196,133 interactants connecting UniProt protein IDs.
@@
[16] consists of 2,656,221 data points involving 1.2 million compounds and 9,000 targets. Rather than using the affinity score, we create triples for every drug-protein pair. To avoid data redundancy, we remove duplicated triples from the TDC DTI dataset. This results in a final set of 2,232,392 triples included in Otter-UBC KG.
@@
[9] consists of drug-like bioactive molecules, in Otter-UBC KG we included 10,261 ChEMBL molecule IDs with their corresponding SMILES downloaded from OpenTargets [18], from which 7,610 have a sameAs link to Drugbank ID molecules.
@@
[13] consists of detailed chemical data on 9,749 drugs (such as SMILES, description, indication, mechanism of action, organism, average mass, toxicity and other calculated and experimental properties), drug pathways and 1,301,422 drug interactions. Due to licenses restrictions, in Otter-UBC we only release drugbank IDs linked from other sources, such as Uniprot and ChemBL.
@@
[17] comprises a collection of 22,886 active compounds and their corresponding affinities towards 102 targets. For our study, we utilized a preprocessed version of the DUDe [31], which includes 1,452,568 instances of drug-target interactions. To prevent any data leakage, we eliminated the negative interactions and the overlapping triples with the TDC DTI dataset. This yielded 40,216 pairs of drug-target interactions.
@@
(the Precision Medicine Knowledge Graph) [4] integrates 20 biomedical resources, it describes 17,080 diseases with 4 million relationships. PrimeKG includes nodes describing Gene/Proteins (29,786) and Drugs (7,957 nodes). Our multimodal KG derived from PrimeKG contains 13 types of modalities and 12,757,300 edges. These edges consist of 154,130 data properties and 12,603,170 object properties. Among them, 642,150 edges represent protein interactions, 25,653 edges represent drug-protein interactions, and 2,672,628 edges represent drug interactions. Our multimodal KG extends the original PrimeKG graph by adding SMILES for listed drugs, and sequences for listed Uniprot proteins.
@@
(Search Tool for Interacting Chemicals) [34] is a database of known and predicted interactions between chemicals represented by SMILES strings and proteins whose sequences are taken from STRING database [33]. For the MKG curation we filtered only the interaction with highest confidence, i.e., the one which is higher 0.9. This resulted into 10,717,791 triples for 17,572 different chemicals and 1,886,496 different proteins. Furthermore, the graph is split into 5 subgraphs of roughly similar size so the GNN is trained sequentially on each of them, building on previous subgraph’s model.